OpenStudy (anonymous):

Calculate the radius of convergence and the interval of convergence (with examination of the endpoints) for the following power series: \(\sum_{n=0}^{\infty} \frac{x^{2n}]{2n+1}\)

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} \frac{x^{2n}}{2n+1}\]

OpenStudy (anonymous):

ratio test should do it for this one

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac{2n+3}{2n+1}=1\] pretty much by your eyeballs so \[|x^2|<1\] and now check at \(x=-1\) and \(x=1\)

OpenStudy (anonymous):

thank you satellite

OpenStudy (anonymous):

yw we still need to check the endpoints

OpenStudy (anonymous):

actually there is very little to check, because at both \(x=1\) and \(x=-1\) you get \[\sum\frac{1}{2n+1}\] which certainly does not converge

OpenStudy (anonymous):

yes i looked again it looks good but difficult to figure out how we came to result

OpenStudy (anonymous):

we can do it more precisely if you like

OpenStudy (anonymous):

yes i would be thankfull for it

OpenStudy (anonymous):